# Descend or Rewind? Stochastic Gradient Descent Unlearning

### v1

@@ 4.3 D2D: Descending with SGD
Finally, in this section, we prove that SGD-D2D achieves certified unlearning. We first show that during training, the biased SGD iterates $\{\theta_{t}\}$ converge to a neighborhood of $\theta^{*^{\prime}}$, the global minimum of $\mathcal{L}_{\mathcal{D}^{\prime}}$, and we then show that the unbiased SGD iterates $\{\theta^{\prime\prime}_{t}\}$ converge closer to $\theta^{*^{\prime}}$. We note that this differs from the original D2D proof, where Lipschitz continuity and strong convexity are combined to show that $\theta^{*^{\prime}}$ is close to the global minimum of $\mathcal{L}_{\mathcal{D}}$. When we exit this highly constrained setting, this property no longer holds.
Theorem 4.
Suppose that the loss function $\ell$ satisfies Assumptions 1, 3, 4, and 5. Let $\eta\leq\frac{1}{BL}$ and $\frac{m}{n}<\frac{1}{6B+1}$. Let
@@
Then SGD-D2D is an $(\epsilon,2\delta)$-certified unlearning algorithm with
$\sigma=\frac{\Sigma}{\epsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}},$
where
@@
Theorem 4 shows that SGD-D2D can achieve certified unlearning as long as the proportion of unlearned data is small enough, which allows the unlearning bias to be “folded into” the standard unbiased SGD analysis, yielding linear convergence to the global minimum. The noise decays exponentially with the number of unlearning iterates $K$. This mirrors the original D2D result in [neel21a], but the stochastic setting yields an $O(\eta)$ constant term at the end which cannot be eliminated. Moreover, without the bounded gradient/Lipschitz smoothness assumption of [neel21a], the computation advantage of unlearning, TKT-K, depends on the initialization; in fact, if $\theta_{0}$ happens to be close enough to the optimum of $\mathcal{L}_{\mathcal{D}^{\prime}}$, then unlearning may not be more efficient than retraining from scratch.
@@ 5 Conclusion
In this work, we prove certified unlearning guarantees for (P)SGD-R2D and SGD-D2D. Our analysis demonstrates that we can achieve a tighter dependence on $\delta$ for D2D-SGD, due to a second moment bound obtained from all trajectories coming close to the same unique minimum. SGD-R2D, which relies on the underlying contractivity of the gradient system, has a looser dependence on $\delta$ because these properties only hold under Riemannian metrics that satisfy the triangle inequality, leading to a first-moment bound. However, SGD-R2D is always more computationally efficient than retraining, and for strongly convex functions it can have a potentially infinite computational advantage.
The next steps include computing utility and generalization guarantees, as well as developing experiments to compare various algorithms. In addition, there are several potential future directions of this work. First, we note that other certified unlearning algorithms, such as [chien2024stochastic] and [koloskova2025certified], follow a D2D-like framework with clipped or noisy gradient updates, instead of vanilla SGD. Combining their approaches with R2D instead could yield improved guarantees, especially on nonconvex functions. Second, our results for PSGD rely on sampling with replacement to achieve tighter bounds. Whenever a sample from $Z$ is selected during the generation of $\theta_{t}$, we can replace it with an i.i.d. sample from $\mathcal{D}^{\prime}$ for producing $\theta^{\prime}_{t}$, maintaining a “valid” run of $\theta^{\prime}_{t}$. However, if sampling occurs without replacement, then drawing one sample affects the probability of the next. Developing the corresponding theory for sampling without replacement, which is more commonly implemented in practice, would be a useful future direction.
Appendix A Additional Discussion of Related Work
Differential Privacy.
Differential privacy (DP) [dwork2006] is a mathematical framework for privacy that leverages noise injection to limit the impact of the inclusion or exclusion of any one data sample on the output of the algorithm.
Definition 3.
[dwork2014algorithmic] A randomized algorithm $\mathcal{M}$ with domain $\mathbb{N}^{|\mathcal{X}|}$ is $(\epsilon,\delta)$-differentially private if for all $S\subset Range(\mathcal{M})$ and for all adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}\in\mathbb{N}^{|\mathcal{X}|}$
$\mathbb{P}[\mathcal{M}(\mathcal{D})\in S]\leq e^{\epsilon}\mathbb{P}[\mathcal{M}(\mathcal{D}^{\prime})\in S]+\delta.$
A common DP technique is the Gaussian mechanism, which scales Gaussian noise to the deterministic global sensitivity, a worst-case bound on how much the algorithm output shifts when the underlying dataset is changed [dwork2014algorithmic]. This technique has been applied to privatizing deterministic functions such as dataset queries or full-batch gradient descent, as well as randomized algorithms such as SGD [abadiDP] [wublackboxdp] [zhangDPijcai2017p548] [chien2025convergent]. To bound the sensitivity, the well-known DP-SGD algorithm [abadiDP] clips the gradient at each step to minimize the impact of any single data sample, and the black-box SGD algorithm [wublackboxdp] requires a uniformly bounded gradient.
Biased SGD.
Our work relies on interpreting the certified unlearning problem as a form of biased or disturbed gradient descent. There are two complementary viewpoints in the literature. First, several works analyze biased SGD from a classic optimization perspective to achieve convergence to a minimum or stationary point. Typically, these works establish minimal viable assumptions on the bias or noise to achieve convergence. For example, [demidovich2023guidezoobiasedsgd] summarizes biased SGD results for nonconvex and strongly convex (or PL) loss functions. In addition, [pmlr-v51-hu16b] considers bandit convex optimization with absolutely bounded bias and on a bounded convex domain, and [driggsbiasedSGD] considers biased SGD schemes that satisfy a specific “memory-biased” property, like SARAH and SVRG. In our work, we utilize the groundwork laid in [AjalloeianBiasedSGD] to show unlearning for D2D on strongly convex functions.
The second viewpoint of biased SGD is through the lens of nonlinear contraction theory [kozachkovgeneralization] [sontag2021remarks], characterizing the bias at every step as disturbances to a noisy gradient system. The trajectory stability, which determines how far a gradient trajectory will diverge when disturbed, can be used to bound the distance between SGD training on two different loss functions. We leverage well-known properties of these systems to analyze stochastic R2D unlearning.
Appendix B Proofs
B.1 Overview
@@
In Appendix B.2, we list some helper lemmas that are standard results in probability and optimization that are useful for this work. In Appendix B.3, we prove the theorems governing the relaxed Gaussian mechanism for first or second moment bounds, which are used for all later proofs. In Appendix B.4, we establish the contractive, semi-contractive, and expansive properties of gradient algorithms on strongly convex, convex, and nonconvex functions respectively. These results, which reflect well-known behaviors of gradient systems, are essential for proving unlearning for PSGD-R2D in Appendix B.5 and SGD-R2D in Appendix B.6. Finally, in Appendix B.7, we prove unlearning for SGD-D2D, under the same unbounded domain conditions of the analysis of SGD-R2D in the previous section.
B.2 Helper lemmas
@@
Lemma 8.
Suppose for all $z\in\mathcal{D}$, $\ell(z;\theta)$ is $L$-Lipschitz smooth for all $\theta\in\mathbb{R}^{d}$. Then $\mathcal{L}_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(z;\theta)$ is also $L$-Lipschitz smooth in $\theta$.
Lemma 9.
@@
Lemma 10.
[karimiPL] Suppose a function $\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the PL inequality (Definition 4) with parameter $\mu>0$. Let $\mathcal{X}=\arg\min_{\theta\in\Theta}\mathcal{L}(\theta)$ represent the solution set of $\mathcal{L}$, and let $\theta^{*}=Proj_{\mathcal{X}}(\theta)=\arg\min_{x\in\mathcal{X}}\lVert x-\theta\rVert$. Then $\mathcal{L}$ satisfies the quadratic growth condition such that
@@
B.3 Relaxed Gaussian mechanism
Lemma 12.
Let $x$ and $y$ be random variables over some domain $\Omega$, and let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$.
  1. 1.
    Suppose the second moment of the distance between $x$ and $y$ is bounded as follows
    $\mathbb{E}[\lVert x-y\rVert^{2}]\leq\Sigma^{2}.$ (11)
    Then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\epsilon,2\delta)$-indistinguishable if
    $\sigma=\frac{\Sigma}{\epsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}}.$
  2. 2.
    Suppose the first moment of the distance between $x$ and $y$ is bounded as follows
    $\mathbb{E}[\lVert x-y\rVert]\leq\Sigma.$ (12)
    Then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\epsilon,2\delta)$-indistinguishable if
    $\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\epsilon\delta}.$
Lemma 13.
@@ B.3 Relaxed Gaussian mechanism
Let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. Then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\epsilon,2\delta)$-indistinguishable if
$\sigma=\frac{\Sigma}{\epsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}}.$
@@
Let $E$ represent the event that $\lVert x-y\rVert\leq\frac{\Sigma}{\sqrt{\delta}}$, such that $\mathbb{P}[E]>1-\delta$ and $\mathbb{P}[\neg E]\leq\delta$. Then when $E$ occurs, the standard Gaussian mechanism holds with sensitivity $\frac{\Sigma}{\sqrt{\delta}}$, and we have for $\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\sqrt{\delta}\epsilon}$
$\mathbb{P}[X\in S|E]\leq e^{\epsilon}\mathbb{P}[Y\in S|E]+\delta.$
@@
$\leq$ $(e^{\epsilon}\mathbb{P}[Y\in S\,|\,E]+\delta)\mathbb{P}[E]+\delta$
$\leq$ $e^{\epsilon}\mathbb{P}[Y\in S\,|\,E]\mathbb{P}[E]+\delta+\delta$
$=$ $e^{\epsilon}\mathbb{P}[E\,|\,Y\in S]\mathbb{P}[Y\in S]+2\delta$
$\leq$ $e^{\epsilon}\mathbb{P}[Y\in S]+2\delta.$
We note that independence is not required for Lemma 13 to hold, as the bounds are true for arbitrary joint distribution between $x$, $y$, $X$ and $Y$.
Lemma 14.
@@
Let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. Then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\epsilon,2\delta)$-indistinguishable if
$\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\epsilon\delta}.$
@@
B.4 Contraction of gradient systems
Lemma 15.
Suppose $\ell$ satisfies Assumption 1. For some data sample $z\sim\mathcal{Z}$, we define the gradient descent map as
@@ B.4 Contraction of gradient systems
B.5 PSGD-R2D
In this section, we prove Theorem 2, which establishes certified unlearning for PSGD-R2D. We define the iterates $\{\theta_{t}\}_{t=0}^{T}$, $\{\theta^{\prime}_{t}\}_{t=0}^{T}$, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$ as follows.
@@ B.5 PSGD-R2D
To prove Theorem 2, we need to determine $\Sigma$ such that $\mathbb{E}[\lVert\theta^{\prime}_{t}-\theta^{\prime\prime}_{K}\rVert]\leq\Sigma$. This result can be combined with the relaxed Gaussian mechanism (Lemma 12) to yield the $(\epsilon,\delta)$ guarantee in Theorem 3. We can determine $\Sigma$ for contracting, semi-contracting, and expanding systems, which correspond to strongly convex, convex, and nonconvex loss functions (Lemma 15). This yields our final noise guarantees in Theorem 2. To bound $\Sigma$, we first require the following Lemma 17 that bounds the expected distance between $\theta^{\prime}_{t}$ and $\theta_{t}$.
Lemma 17.
Suppose that the loss function $\ell$ satisfies Assumptions 1 and within the closed, bounded, and convex set $\mathcal{C}$, the gradient of $\ell$ is uniformly bounded by some constant $G\geq 0$ such that for all $z\in\mathcal{Z}$ and $\theta\in\mathcal{C}$, $\lVert\nabla\ell(z;\theta)\rVert\leq G.$ Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the following property for some $\gamma>0$,
@@
Lemma 17 allows us to prove Theorem 18, which can be combined with Lemma 15 and Lemma 12 to yield the result in Theorem 2.
Theorem 18.
Suppose that the loss function $\ell$ satisfies Assumptions 1 and within the closed, bounded, and convex set $\mathcal{C}$, the gradient of $\ell$ is uniformly bounded by some constant $G\geq 0$ such that for all $z\in\mathcal{Z}$ and $\theta\in\mathcal{C}$, $\lVert\nabla\ell(z;\theta)\rVert\leq G.$ Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the following property for some $\gamma>0$,
@@
Lemma 17 shows that the expected distance between $\theta_{t}$ and $\theta^{\prime}_{t}$ is bounded as a function of $t$. So for $\gamma\neq 1$, we have
@@
From Lemma 16, we have that we can choose a coupling of $\{\theta^{\prime}_{t}\}_{t=T-K}^{T}$, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$, such that
@@
Theorem 2 follows directly from Theorem 18, Lemma 12 and Lemma 15.
B.5.1 Proof of Corollary 2.1
@@
B.6 SGD-R2D
In this section, we prove Theorem 3, which establishes certified unlearning for SGD-R2D without projection or bounded domain. We establish preliminaries and present Lemmas 24 and 20. We then use those results to prove Theorem 22, from which Theorem 3 directly follows. In Appendix B.7.1, we prove Lemma 24. In Appendix B.6.2, we prove Lemma B.6.2.

### v2

@@ 4.3 D2D: Descending with SGD
Finally, we prove that SGD-D2D achieves certified unlearning. We first show that during training, the biased SGD iterates $\{\theta_{t}\}$ converge to a neighborhood of $\theta^{*^{\prime}}$, the global minimum of $\mathcal{L}_{\mathcal{D}^{\prime}}$, and we then show that the unbiased SGD iterates $\{\theta^{\prime\prime}_{t}\}$ converge closer to $\theta^{*^{\prime}}$. We note that this differs from the original D2D proof, where Lipschitz continuity and strong convexity are combined to show that $\theta^{*^{\prime}}$ is close to the global minimum of $\mathcal{L}_{\mathcal{D}}$, a property that does not hold outside of this highly constrained setting.
Theorem 4.10.
Suppose that $\ell$ satisfies Assumptions 4.1, 4.3, 4.7, and 4.8. Let $\eta\leq\frac{1}{BL}$ and $\frac{m}{n}<\frac{1}{6B+1}$. Let
@@
Then for $0<\varepsilon\leq 1$ and $\delta>0$, SGD-D2D (Algorithms 3 and 4) is an $(\varepsilon,2\delta)$-certified unlearning algorithm with $\sigma=\frac{\Sigma}{\varepsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}},$ where
@@
Theorem 4.10 shows that SGD-D2D can achieve certified unlearning as long as the proportion of unlearned data is small enough, which allows the unlearning bias to be “folded into” the standard unbiased SGD analysis, yielding linear convergence to the global minimum. The noise decays exponentially with the number of unlearning iterates $K$. This mirrors the original D2D result in (Neel et al., 2021), but the stochastic setting yields an $O(\eta)$ constant term at the end which cannot be eliminated. Moreover, without the bounded gradient/Lipschitz smoothness assumption of (Neel et al., 2021), the computation advantage of unlearning, TKT-K, depends on the initialization; in fact, if $\theta_{0}$ happens to be close enough to the optimum of $\mathcal{L}_{\mathcal{D}^{\prime}}$, then unlearning may not be more efficient than retraining from scratch.
5 Experiments
Figure 1: Comparison of unlearning capabilities of noiseless R2D and D2D. The eICU results are in the top row and the Lacuna-100 results are in the bottom row. The left two columns display the L2 distance from unlearning and the original model as the number of unlearning iterations increases. The right two columns display the model performance on the unlearned, retained, and test sets, denoted as $\mathcal{D}_{unlearn}$, $\mathcal{D}_{train}$, and $\mathcal{D}_{test}$ respectively.
We follow the experimental setup in (Mu and Klabjan, 2025) applied to PSGD instead of GD. For all experiments, we train a binary classifier with SGD and the cross-entropy loss function, and we unlearn a subset of the data from the trained model. For small-scale experiments, we use the eICU dataset (Pollard et al., 2018), a large multi-center intensive care unit (ICU) tabular dataset, and we train a multilayer perceptron (MLP) with three hidden layers to predict whether patients stay in the hospital for longer or shorter than a week, using their intake data. For large-sclae experiments, we consider the Lacuna-100 dataset (Golatkar et al., 2019), constructed from the VGGFace2 dataset (Cao et al., 2018) and the MAAD-Face annotations (Terhörst et al., 2020) as described in (Mu and Klabjan, 2025), and we train a ResNet-18 model (He et al., 2016) to perform binary gender classification. In contrast to prior work, our experiments focus on examining the difference between descending vs. rewinding. We considering varying $K$, or number of unlearning iterations, and we evaluate the effects of D2D and R2D at the same level of computation and batch size. For additional experimental details, see Appendix B.
We evaluate unlearning efficacy using three metrics. First, we compute the L2 distance in parameter space between unlearning and retraining, where better unlearning minimizes this distance. Second, we consider the model performance on the unlearned dataset $\mathcal{D}_{unlearn}$ before and after unlearning. A decrease in performance suggests that the model is losing information about the unlearned samples. Finally, we apply membership inference attacks (MIAs) which attempt to distinguish between unlearned data and data that has never been in the training dataset. A less successful attack, indicated by a lower Area Under the Receiver Operating Characteristic Curve (AUC), indicates that unlearning is more successful. We utilize both the classic MIA (Shokri et al., 2017) that considers the model outputs after unlearning as well as the more sophisticated unlearning MIA (MIA-U) (Chen et al., 2021) which compares the model outputs before and after unlearning.
Figure 1 compares the first two unlearning metrics, L2 distance and model performance, for R2D and D2D. Our two experimental settings illustrate varying practical outcomes of the unlearning methods. For eICU, R2D has a reliable outcome: the unlearned model moves away from the original model and towards the retrained model. This is reflected in the L2 distance as well as the model performance on $\mathcal{D}_{unlearn}$, which decreases with increased rewinding even as the performance on the retained and test sets remain constant. In contrast, D2D causes the model to progress away from both the original and retrained models. The model performance on all data, including $\mathcal{D}_{unlearn}$, improves, suggesting that the optimization algorithm has found a new descent direction on the new loss function $\mathcal{L}_{\mathcal{D}^{\prime}}$. For Lacuna-100, rewinding still reliably moves the model away from the original trained model, but the L2 distance from the retrained model does not change significantly. This may be because in the stochastic setting, R2D is only theoretically guaranteed to reduce the distance in expectation. However, R2D does still reduce the model performance on $\mathcal{D}_{unlearn}$ while keeping the performance on other sets constant. In contrast, D2D displays very little movement in parameter space or in model performance, suggesting that it is stalled at a stationary point at initialization. Moreover, a large number of descent steps risks overfitting, as shown in the decrease in performance on $\mathcal{D}_{test}$.
Figure 2 displays the results of the MIAs. While the output of the MIAs can be highly nonlinear (especially for non-i.i.d. data), in general we observe that more unlearning decreases the attack success, especially for rewinding. Moreover, adding a small amount of noise tends to reduce the success of the attack. Ultimately, we find that descent-based algorithms (including D2D, (Chien et al., 2024a), and (Koloskova et al., 2023)) can either improve the model performance on all sets, which may obfuscate whether unlearning has occurred, or it may stall at a stationary point. On the other hand, R2D has a more reliable unlearning effect, but may not get any performance boost from unlearning.
For additional results and experimental details, including the privacy-utility-complexity tradeoff of the derived guarantees, comparisons with baseline certified unlearning methods, and anonymous GitHub repository, see Appendix B.
Figure 2: Comparison of MIA success for R2D and D2D. The left column displays the results of the classic MIA, and the right column displays the results of the unlearning MIA.
@@ 6 Conclusion
In this work, we prove certified unlearning guarantees for SGD-R2D and SGD-D2D. Our analysis demonstrates that we can achieve a tighter bound for SGD-D2D on strongly convex functions, whereas SGD-R2D, which relies on the underlying contractivity of the gradient system, has a looser first-moment bound. However, SGD-R2D is always more computationally efficient than retraining. Experiments demonstrate the practical effects of each method.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs
A.1 Overview
@@
In Appendix A.2, we list some helper lemmas that are standard results in probability and optimization that are useful for this work. In Appendix A.3, we prove the theorems governing the first or second moment sensitivity bounds, which are used for all later proofs. In Appendix A.4, we establish the contractive, semi-contractive, and expansive properties of gradient algorithms on strongly convex, convex, and nonconvex functions respectively. These results, which reflect well-known behaviors of gradient systems, are essential for proving unlearning for PSGD-R2D in Appendix A.5 and SGD-R2D in Appendix A.6. Finally, in Appendix A.7, we prove unlearning for SGD-D2D, under the same unbounded domain conditions of the analysis of SGD-R2D in the previous section.
A.2 Helper lemmas
@@
Lemma A.4.
Suppose for all $z\in\mathcal{D}$, $\ell(z;\theta)$ is $L$-Lipschitz smooth for all $\theta\in\mathbb{R}^{d}$. Then $\mathcal{L}_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(z;\theta)$ is aSlso $L$-Lipschitz smooth in $\theta$.
Lemma A.5.
@@
Lemma A.7.
(Karimi et al., 2016) Suppose a function $\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the PL inequality (Definition A.6) with parameter $\mu>0$. Let $\mathcal{X}=\arg\min_{\theta\in\Theta}\mathcal{L}(\theta)$ represent the solution set of $\mathcal{L}$, and let $\theta^{*}=Proj_{\mathcal{X}}(\theta)=\arg\min_{x\in\mathcal{X}}\lVert x-\theta\rVert$. Then $\mathcal{L}$ satisfies the quadratic growth condition such that
@@
A.3 Proof of Lemma 4.4
Lemma A.9.
@@
Let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. Then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\varepsilon,2\delta)$-indistinguishable if
$\sigma=\frac{\Sigma}{\varepsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}}.$
@@
Let $E$ represent the event that $\lVert x-y\rVert\leq\frac{\Sigma}{\sqrt{\delta}}$, such that $\mathbb{P}[E]>1-\delta$ and $\mathbb{P}[\neg E]\leq\delta$. Then when $E$ occurs, the standard Gaussian mechanism holds with sensitivity $\frac{\Sigma}{\sqrt{\delta}}$, and we have for $\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\sqrt{\delta}\varepsilon}$
$\mathbb{P}[X\in S|E]\leq e^{\varepsilon}\mathbb{P}[Y\in S|E]+\delta.$
@@
$\leq$ $(e^{\varepsilon}\mathbb{P}[Y\in S\,|\,E]+\delta)\mathbb{P}[E]+\delta$
$\leq$ $e^{\varepsilon}\mathbb{P}[Y\in S\,|\,E]\mathbb{P}[E]+\delta+\delta$
$=$ $e^{\varepsilon}\mathbb{P}[E\,|\,Y\in S]\mathbb{P}[Y\in S]+2\delta$
$\leq$ $e^{\varepsilon}\mathbb{P}[Y\in S]+2\delta.$
We note that independence is not required for Lemma A.9 to hold, as the bounds are true for arbitrary joint distribution between $x$, $y$, $X$ and $Y$.
Lemma A.10.
@@
Let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. Then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\varepsilon,2\delta)$-indistinguishable if
$\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\varepsilon\delta}.$
@@
A.4 Contraction of gradient systems
Lemma A.11.
Suppose $\ell$ satisfies Assumption 4.1. For some data sample $z\sim\mathcal{Z}$, we define the gradient descent map as
@@ A.4 Contraction of gradient systems
A.5 PSGD-R2D
In this section, we prove Theorem 4.5, which establishes certified unlearning for PSGD-R2D. We define the iterates $\{\theta_{t}\}_{t=0}^{T}$, $\{\theta^{\prime}_{t}\}_{t=0}^{T}$, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$ as follows.
@@ A.5 PSGD-R2D
To prove Theorem 4.5, we need to determine $\Sigma$ such that $\mathbb{E}[\lVert\theta^{\prime}_{t}-\theta^{\prime\prime}_{K}\rVert]\leq\Sigma$. This result can be combined with Lemma 4.4 to yield the $(\varepsilon,\delta)$ guarantee in Theorem 4.9. We can determine $\Sigma$ for contracting, semi-contracting, and expanding systems, which correspond to strongly convex, convex, and nonconvex loss functions (Lemma A.11). This yields our final noise guarantees in Theorem 4.5. To bound $\Sigma$, we first require the following Lemma A.13 that bounds the expected distance between $\theta^{\prime}_{t}$ and $\theta_{t}$.
Lemma A.13.
Suppose that the loss function $\ell$ satisfies Assumptions 4.1 and within the closed, bounded, and convex set $\mathcal{C}$, the gradient of $\ell$ is uniformly bounded by some constant $G\geq 0$ such that for all $z\in\mathcal{Z}$ and $\theta\in\mathcal{C}$, $\lVert\nabla\ell(z;\theta)\rVert\leq G.$ Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the following property for some $\gamma>0$,
@@
Lemma A.13 allows us to prove Theorem A.14, which can be combined with Lemma A.11 and Lemma 4.4 to yield the result in Theorem 4.5.
Theorem A.14.
Suppose that the loss function $\ell$ satisfies Assumptions 4.1 and within the closed, bounded, and convex set $\mathcal{C}$, the gradient of $\ell$ is uniformly bounded by some constant $G\geq 0$ such that for all $z\in\mathcal{Z}$ and $\theta\in\mathcal{C}$, $\lVert\nabla\ell(z;\theta)\rVert\leq G.$ Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the following property for some $\gamma>0$,
@@
Lemma A.13 shows that the expected distance between $\theta_{t}$ and $\theta^{\prime}_{t}$ is bounded as a function of $t$. So for $\gamma\neq 1$, we have
@@
From Lemma A.12, we have that we can choose a coupling of $\{\theta^{\prime}_{t}\}_{t=T-K}^{T}$, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$, such that
@@
Theorem 4.5 follows directly from Theorem A.14, Lemma 4.4 and Lemma A.11.
A.5.1 Proof of Corollary 4.6
@@
A.6 SGD-R2D
In this section, we prove Theorem 4.9, which establishes certified unlearning for SGD-R2D without projection or bounded domain. We establish preliminaries and present Lemmas A.20 and A.16. We then use those results to prove Theorem A.18, from which Theorem 4.9 directly follows. In Appendix A.7.1, we prove Lemma A.20. In Appendix A.6.2, we prove Lemma A.6.2.